Lloyd Inc. has sales of $450,000, a net income of $45,000, and the following balance sheet: Cash $67,500 Accounts payable $60,075 Receivables 124,200 Notes payable to bank 39,150 Inventories 330,750 Total current liabilities $99,225 Total current assets $522,450 Long-term debt 84,375 Net fixed assets 152,550 Common equity 491,400 Total assets $675,000 Total liabilities and equity $675,000 The new owner thinks that inventories are excessive and can be lowered to the point where the current ratio is equal to the industry average, 2.5x, without affecting sales or net income. If inventories are sold and not replaced (thus reducing the current ratio to 2.5x); if the funds generated are used to reduce common equity (stock can be repurchased at book value); and if no other changes occur, by how much will the ROE change? Do not round intermediate calculations. Round your answer to two decimal places.

What will be the firm's new quick ratio? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

To make current ratio = 2.5, current assets will be equal to Current Liabilities*2.5

= 99,225*2.5

= $248,062.5

Decrease in Current Assets = 522,450-248,062.5

= $274,387.5

Equity will reduce by the same amount

ROE = Net Income/Equity

ROE earlier = 45,000/491,400 = 9.16%

New ROE = 45,000/217,012.5 = 20.74%

Hence, ROE will change by 20.74%-9.16% = 11.58%

Quick Ratio = (Current Assets - Inventory)/Current Liabilities

= 191,700/99,225

= 1.93:1